OpenStudy (softballgirl372015):

Can someone please explain how to graph the function y=tan(x-pi/6). I am confused with the phase shift involving pi.

jimthompson5910 (jim_thompson5910):

because the x term has a coefficient of 1, you are shifting all of tan(x) pi/6 units to the right

jimthompson5910 (jim_thompson5910):

In general, if you have y = A*tan(Bx-C) + D the phase shift is C/B and that tells you how to shift the graph right or left

OpenStudy (softballgirl372015):

So at what points would the starting and ending values of this tan graph be?

jimthompson5910 (jim_thompson5910):

not sure what you mean

jimthompson5910 (jim_thompson5910):

the graph of tan(x) goes on forever in both directions along the x axis

OpenStudy (softballgirl372015):

I guess a better way to phrase the question is at what points would the asymptotes be for this graph?

jimthompson5910 (jim_thompson5910):

well the asymptotes for tan(x) are at x = (pi/2)*n where n is any integer

jimthompson5910 (jim_thompson5910):

this is because tan(x) = sin(x)/cos(x) and cos(x) = 0 when x = (pi/2)*n (n is any integer)

jimthompson5910 (jim_thompson5910):

so you would shift everything over (including the asymptotes) pi/6 units to the right to get y=tan(x-pi/6).

OpenStudy (softballgirl372015):

Okay I drew it out and I understand it now : )

jimthompson5910 (jim_thompson5910):

ok that's great